Integration by Parts Fractions


1.) ∫(8t^3 + 13)dt/(t+2)(4t^2+1)


2.) ∫(5x^2 -3x+ 18)dx/x(9-x^2) from 1 to 2


1
Expert's answer
2021-05-07T09:51:50-0400

"\displaystyle
(1)\\
\frac{8t^3 + 13}{(t+2)(4t^2+1)}= 2 - \frac{3}{2 + t} - {4t - 6}{1 + 4 t^2}\\

\begin{aligned}
\int \frac{8t^3 + 13}{(t+2)(4t^2+1)}\,\mathrm{d}t &= \int 2 - \frac{3}{2 + t} - \frac{4t - 6}{1 + 4 t^2}\,\mathrm{d}t 
\\& = 2t - 3\ln(2 + t)
\\& - \frac{\ln(1 + 4t^2)}{2} + 3\arctan(2t) + C
\end{aligned}\\

(2)\\
\frac{5x^2 -3x+ 18}{x(9-x^2)} = \frac{2}{x} - \frac{4}{x + 3} - \frac{3}{x - 3}\\

\begin{aligned}
\int\frac{5x^2 -3x+ 18}{x(9-x^2)} \,\mathrm{d}x &= \int\frac{2}{x}\mathrm{d}x - \int\frac{4}{x + 3} + \int \frac{3}{3 - x}\mathrm{d}x
\\&= 2\ln{x} - 4\ln(x + 3) - 3\ln(3 - x) + C\\
\int_1^2\frac{5x^2 -3x+ 18}{x(9-x^2)} \,\mathrm{d}x &= 2\ln(2) - 4\ln(5) + 4\ln(4) + 3\ln(2)
\\& = 13\ln(2) - 4\ln(5)
\end{aligned}"